# KA-GNN: Kolmogorov-Arnold Graph Neural Networks for Molecular Property Prediction

### v1

@@ 3.4 KA-GNN Method
$\begin{split}h_{v}^{(0)}&:=\left(f_{v}\oplus(\frac{1}{|N(v)|}\sum_{u\in N(v)}f_{(u,v)})\right),\\ h_{uv}^{(0)}&:=\text{Linear}\left(f_{v}\right)+\text{Linear}\left(f_{uv}\right)+\text{Linear}\left(f_{u}\right)\end{split}$ (7)
During the message passing process of the Mol-GAT, we use the initial features to calculate the attention score for the aggregation operation.
@@
$\begin{split}z_{v}^{(l+1)}&:=\text{Linear}\left(h_{v}^{(l)}\right),z_{u}^{(l+1)}:=\text{Linear}\left(h_{u}^{(l)}\right),\\ \alpha_{uv}&:=\text{Softmax}\left(h_{uv}^{(l)}\right),m_{u}^{(l+1)}:=z_{u}^{(l+1)}\cdot\alpha_{uv}\\ m_{v}^{(l+1)}&:=z_{v}^{(l+1)}+\left(\sum_{u\in N(v)}m_{u}^{(l+1)}\right)\end{split}$ (8)
Then we use KAN for information updating.
$h_{v}^{(l+1)}:=\text{KAN}\left(m_{v}^{(l+1)}\right),h_{uv}^{(l+1)}:=\text{KAN}(h_{uv}^{(l)})$ (9)
Readout
Following the message passing process, we apply average pooling over all node feature vectors to obtain a final representation of the molecular graph $G$. Subsequently, we apply a one-layer or two-layer KAN as a readout function on this representation for the prediction task.
The loss function employed in our model is the cross-entropy loss function, defined as follows:
$\mathcal{L}=-\sum_{i}\left(y_{i}\log(\hat{y}_{i})+(1-y_{i})\log(1-\hat{y}_{i})\right),$
where $y_{i}$ denotes the actual label, and $\hat{y}_{i}$ denotes the predicted label. The architecture of the KA-GNN and KA-GAT model is illustrated in Figure 1.
4 Experiments
Table 2: Performance metrics of various models across different datasets, The best performance is indicated as bold and the second-best performance is in bold. Standard deviations are indicated as subscripts.
@@ 4 Experiments
D-MPNN 0.809(0.006) 0.710(0.003) 0.906(0.007) 0.570(0.007) 0.759(0.007) 0.771(0.005) 0.786(0.014)
AttentiveFP 0.784(0.022) 0.663(0.018) 0.847(0.003) 0.606(0.032) 0.781(0.005) 0.757(0.014) 0.786(0.015)
N-GramRF 0.779(0.015) 0.697(0.006) 0.775(0.040) 0.668(0.007) 0.743(0.009) 0.772(0.004) 0.769(0.002)
N-GramXGB 0.791(0.013) 0.691(0.008) 0.875(0.027) 0.655(0.007) 0.758(0.009) 0.787(0.004) 0.748(0.002)
PretrainGNN 0.845(0.007) 0.687(0.013) 0.726(0.015) 0.627(0.008) 0.781(0.006) 0.799(0.007) 0.813(0.021)
GROVE_base 0.821(0.007) 0.700(0.001) 0.812(0.030) 0.648(0.006) 0.743(0.001) 0.625(0.009) 0.673(0.018)
GROVE_large 0.810(0.014) 0.695(0.001) 0.762(0.037) 0.654(0.001) 0.735(0.001) 0.682(0.011) 0.673(0.018)
GraphMVP 0.812 (0.009) 0.724 (0.016) 0.791 (0.028) 0.639 (0.012) 0.759 (0.005) 0.770 (0.012) 0.777 (0.006)
MolCLR 0.824 (0.009) 0.722 (0.021) 0.912 (0.035) 0.589 (0.014) 0.750 (0.002) 0.781 (0.005) 0.796 (0.019)
GEM 0.856(0.011) 0.724(0.004) 0.901(0.013) 0.672(0.004) 0.781(0.001) 0.806(0.009) 0.817(0.005)
Mol-GDL 0.863(0.019) 0.728(0.019) 0.966(0.002) 0.831(0.002) 0.794(0.005) 0.808(0.007) 0.675(0.014)
Uni-mol 0.857 (0.002) 0.729 (0.006) 0.919 (0.018) 0.659 (0.013) 0.796 (0.005) 0.808 (0.003) 0.821 (0.013)
SMPT 0.873 (0.015) 0.734 (0.003) 0.927 (0.002) 0.676 (0.050) 0.797 (0.001) 0.812 (0.001) 0.822 (0.008)
KA-GNN 0.890(0.014) 0.787(0.014) 0.989(0.003) 0.842(0.001) 0.799(0.005) 0.821(0.005) 0.834(0.009)
KA-GAT 0.882(0.007) 0.770(0.021) 0.994(0.003) 0.846(0.002) 0.800(0.004) 0.815(0.008) 0.825(0.005)
4.1 Dataset
4.2 Baselines
Our KA-GNN model has been evaluated alongside a selection of state-of-the-art GNN architectures, including both pre-trained and non-pre-trained models. These encompass MPNN frameworks like D-MPNN [44], attention-driven models such as AttentiveFP [45], and multi-scale approaches exemplified by Mol-GDL [27]. Importantly, our comparisons also extend to geometry-focused graph models that leverage pre-training, including N-Gram [51], PretrainGNN [52], GROVER [25], GraphMVP [66], MolCLR [54], GEM [47], Uni-mol [55], and SMPT [56].
4.3 Setting
In our study, to specifically enhance the model’s performance across various datasets, we selected appropriate hyperparameter settings (such as learning rate, batch size, etc.), details of which can be found in Supplementary Materials. We set the epochs to 1000 and executed the code on a machine equipped with a NVIDIA RTX A5000 32GB GPUs. Each dataset was evaluated five times, calculating the mean as the final metric and recording the standard deviation.
Figure 2: Powerful Fourier KAN fits Polynomial function by adjusting the parameter $K$.
4.4 Results
The comparative results in Table 2 further confirm the superiority of the KA-GNN model. Our model achieves state-of-the-art performance across all benchmark datasets, excelling particularly on complex and challenging datasets such as ClinTox and MUV. These results demonstrate the robust capability of our model in handling molecular data.
We attribute the effectiveness and power of the KA-GNN model to three core innovations:
First, in constructing molecular graphs, we incorporate edges not only from traditional covalent bonds but also from non-covalent interactions based on a cut-off distance. This approach significantly enhances the model’s understanding of molecular structures, enabling it to capture more molecular properties than traditional covalent interactions alone.
Second, we integrate the Kolmogorov-Arnold Network (KAN) into our model. KAN significantly reduces the number of parameters while providing high interpretability and enhanced expressive power, leading to superior performance of the KA-GNN model.
@@ 4.4 Results
Third, we replace the pre-activation function in the KAN model from the original B-splines to Fourier series functions. Fourier series functions, widely used in signal processing, have proven effective in our model for analyzing and learning high-dimensional graphical data. By learning and optimizing the coefficients of these Fourier series functions for different variables, KA-GNN improves the accuracy and stability of predictions when dealing with complex chemical and biological data structures.
These three innovations collectively contribute to the outstanding performance of KA-GNN, making it a powerful tool for molecular property prediction.
Powerful Fourier KAN
The parameter $K$, representing the number of harmonics, is pivotal in the Fourier KAN model. Increasing $K$ allows the model to include a greater number of Fourier series terms in the approximation, thereby enhancing its capacity to approximate more complex functions. However, increasing $K$ also amplifies computational costs and the risk of overfitting. Conversely, smaller $K$ values mitigate these risks but may compromise the model’s expressiveness. Thus, selecting an optimal $K$ is critical for balancing expressiveness with computational efficiency in the Fourier KAN model. Figure 2 shows the fitting of different $K$ values on Polynomial function. More details for Fourier KAN can been found in Appendix A.4.
Ablation Study
Our model utilizes Fourier-based Kolmogorov-Arnold Network (KAN) as the underlying framework, which is very different from traditional MLP-based GNNs. To evaluate the performance of this GNN architecture in predicting molecular properties, we tested GNN models incorporating MLP, original KAN, and Fourier-based KAN. These models are denoted as GNN(MLP), GNN(KAN), KA-GNN, GAT(MLP), GAT(KAN) and KA-GAT respectively. Table 3 summarizes the results, which clearly demonstrate that the KA-GNN model significantly outperforms both the GNN(MLP) and GNN(KAN) models, delivering superior performance across various benchmark datasets.
Table 3: Comparison of the performance of GNN and GAT models across various datasets.
Dataset GNN Models GAT Models
GNN(MLP) GNN(KAN) KA-GNN GAT(MLP) GAT(KAN) KA-GAT
BACE 0.835(0.014) 0.771(0.012) 0.890(0.014) 0.834(0.012) 0.808(0.009) 0.882(0.007)
BBBP 0.735(0.011) 0.723(0.008) 0.787(0.014) 0.707(0.007) 0.657(0.004) 0.770(0.021)
ChinTox 0.979(0.004) 0.973(0.006) 0.989(0.003) 0.983(0.006) 0.948(0.003) 0.994(0.002)
SIDER 0.834(0.001) 0.824(0.003) 0.842(0.001) 0.836(0.002) 0.825(0.004) 0.845(0.002)
Tox21 0.747(0.006) 0.724(0.005) 0.799(0.005) 0.751(0.007) 0.731(0.012) 0.800(0.004)
HIV 0.762(0.005) 0.753(0.007) 0.821(0.005) 0.761(0.003) 0.744(0.018) 0.815(0.008)
MUV 0.741(0.006) 0.638(0.008) 0.834(0.009) 0.784(0.019) 0.792(0.013) 0.825(0.005)
5 Conclusion
The KA-GNN model represents a significant advancement in molecular characterization by leveraging molecular dynamics to construct graphs that incorporate both covalent and non-covalent bonds. Built upon the KAN architecture with Fourier series functions as pre-activation functions, this novel graph neural network has been rigorously evaluated using the MoleculeNet dataset for various downstream tasks. The KA-GNN model consistently outperforms established baselines, demonstrating superior predictive performance.
Ablation studies underscore the model’s value, highlighting its significant contributions to the field of molecular property prediction. Additionally, the integration of Fourier series functions ensures that the KAN model remains interpretable through effective visualization techniques. In summary, the KA-GNN model not only achieves excellent performance but also does so efficiently, without requiring extensive computational resources, making it a valuable tool for advancing molecular research.
Appendix A Appendix / supplemental material
A.1 Main Proof
@@
A.2 Setting
In our study, to specifically enhance the model’s performance across various datasets, we selected appropriate hyperparameter settings (such as learning rate, batch size, etc.), details of which can be found in Table 4. These parameters were chosen based on preliminary experimental evaluations to ensure optimal processing conditions for each dataset. We set the epochs to 1000 and executed the code on a machine equipped with a NVIDIA RTX A5000 32GB GPUs. Each dataset was evaluated five times, calculating the mean as the final metric and recording the standard deviation.
Table 4: Configuration parameters for various datasets
Dataset Batch size LR K #Layers #Parameters
BACE 128 1e-4 1 3 44K
BBBP 128 1e-4 2 1 54K
ChinTox 128 1e-4 2 2 62K
SIDER 128 1e-4 2 1 64K
Tox21 512 1e-4 2 2 63K
HIV 512 1e-4 2 2 62K
MUV 512 1e-4 2 2 64K
A.3 Sensitivity
In this section, we study the effects of harmonic parameter $K$, cutoff distance and model hyperparameters on model robustness.
Number of Harmonics $K$
In optimizing the KAN model using Fourier series functions, the number of harmonics $K$ plays a crucial role in determining its accuracy. A higher $K$ allows the Fourier series to include more high-frequency components, enabling a more detailed representation of complex structures. However, selecting an excessively high $K$ can lead to overfitting, especially when dealing with noisy data, as the series may fit the noise too closely. Table 5 illustrates the impact of different $K$ values on the performance of the KA-GNN model. The results indicate that $K=2$ is the optimal choice for most benchmark datasets.
Table 5: Configuration parameters K for various datasets
Dataset 1 2 3 4 5
BACE 0.890 0.780 0.660 0.715 0.721
BBBP 0.735 0.787 0.766 0.665 0.662
ChinTox 0.989 0.989 0.985 0.981 0.978
SIDER 0.840 0.842 0.842 0.835 0.830
Tox21 0.784 0.799 0.764 0.756 0.741
HIV 0.796 0.821 0.780 0.771 0.749
MUV 0.783 0.834 0.696 0.678 0.643
Cut-off distance
The KA-GNN model integrates both chemical and non-chemical bonds in constructing the molecular graph, thereby leveraging a broader spectrum of molecular information for predicting molecular properties. To evaluate the impact of non-chemical bonds, we introduced a cut-off threshold to select these bonds. As shown in Table 6, when the cut-off distance is set to 0Å, meaning the model does not include any non-chemical bonds identified by the cut-off method, its predictive performance does not show any advantage over scenarios where the cutoff is greater than 0Å. This underscores the importance of non-chemical bonds in accurately representing molecular structures. Notably, setting the cut-off distance to 5Å yields the best predictive results across all benchmark datasets.
Table 6: Cutoff for various datasets
Dataset 0 1 2 3 4 5
BACE 0.839 0.838 0.872 0.860 0.862 0.890
BBBP 0.715 0.709 0.713 0.715 0.767 0.787
ChinTox 0.975 0.977 0.983 0.985 0.982 0.989
SIDER 0.835 0.836 0.835 0.838 0.836 0.842
Tox21 0.748 0.732 0.733 0.763 0.756 0.799
HIV 0.780 0.767 0.774 0.767 0.799 0.821
MUV 0.712 0.709 0.727 0.695 0.749 0.834
Hyperparameters
Table 7 explores the impact of various hyper parameters, including batch size, learning rate (LR), and the number of network layers, on the performance of molecular property prediction models. The results indicate that choosing an appropriate batch size and a lower learning rate typically enhances model performance. Conversely, increasing the number of layers does not always result in performance improvements and can sometimes lead to over-fitting or diminished returns. These observations underscore the importance of careful hyper parameter tuning to achieve optimal model performance.
Table 7: Hyperparameters for various datasets

### v2

@@ 2.2.2 KA-GAT Model
$\begin{split}h_{v}^{(0)}&:=\text{KAN}\left(f_{v}\oplus\left(\frac{1}{|N(v)|}\sum_{u\in N(v)}f_{uv}\right)\right),\\ h_{uv}^{(0)}&:=\text{KAN}\left(\text{Linear}(f_{v})+\text{Linear}(f_{uv})+\text{Linear}(f_{u})\right)\end{split}$ (7)
Subsequently, the KA-GAT model uses these initial embeddings to compute attention scores that guide the aggregation of messages:
@@
$\begin{split}z_{v}^{(l+1)}&:=\text{Linear}\left(h_{v}^{(l)}\right),z_{u}^{(l+1)}:=\text{Linear}\left(h_{u}^{(l)}\right),\\ \alpha_{uv}&:=\text{Softmax}\left(h_{uv}^{(l)}\right),m_{uv}^{(l+1)}:=z_{u}^{(l+1)}\cdot\alpha_{uv}\\ m_{v}^{(l+1)}&:=z_{v}^{(l+1)}+\frac{1}{|N(v)|}\left(\sum_{u\in N(v)}m_{uv}^{(l+1)}\right)\end{split}$ (8)
Then, one or more KAN layers are used to update the node and edge features:
$h_{v}^{(l+1)}:=\text{KAN}(m_{v}^{(l+1)}),\quad h_{uv}^{(l+1)}:=\text{KAN}(h_{uv}^{(l)})$ (9)
Finally, after message passing, the KA-GAT model employs the same readout module as used in the KA-GCN.
2.3 KA-GNNs for Molecular Property Analysis
2.3.1 Performance and comparison of KA-GNNs
To evaluate the performance of our proposed KA-GNN models, we consider seven widely-used benchmark datasets from MoleculeNet [47]. Of these, three datasets—MUV, HIV, and BACE—are from the biophysics domain. The rest four datasets are from the physiology domain, including BBBP, Tox21, SIDER, and ClinTox.
In our comparative analysis across seven datasets, we evaluated a range of geometric deep learning models, including AttentiveFP [48], D-MPNN [49], Mol-GDL [28], N-Gram [50], PretrainGNN [51], GROVER [26], GraphMVP [52], MolCLR [53], GEM [54], Uni-mol [55], SMPT [56], and GNN-SKAN [38]. Additionally, we included our KAN-based GNN models, such as the GNN-SKAN [38]. Further details on these models can be found in Subsection Baseline Models.
The comparative results in Table 2 further confirm the superiority of the KA-GNNs model. Our model achieves state-of-the-art performance across all benchmark datasets, particularly excelling on complex and challenging datasets such as ClinTox and MUV. These results demonstrate the robust capability of our model in handling molecular data. Notably, in the BBBP dataset, the AUC for KA-GCN and KA-GAT showed improvements of approximately 7.95% and 7.68%.
Table 2: Comparison of KA-GNNs with state-of-the-art geometric deep learning models in different molecular property datasets. The performance metric is the average of Receiver Operating Characteristic - Area Under the Curve (ROC-AUC). The standard deviation results are denoted as subscripts. The best model for each category is highlighted in bold, while the second-best performance is marked with underline. The notation - indicates that the results were not reported.
@@ 2.3.1 Performance and comparison of KA-GNNs
D-MPNN [49] 0.809(0.006) 0.710(0.003) 0.906(0.007) 0.570(0.007) 0.759(0.007) 0.771(0.005) 0.786(0.014)
AttentiveFP [48] 0.784(0.022) 0.663(0.018) 0.847(0.003) 0.606(0.032) 0.781(0.005) 0.757(0.014) 0.786(0.015)
N-GramRF [50] 0.779(0.015) 0.697(0.006) 0.775(0.040) 0.668(0.007) 0.743(0.009) 0.772(0.004) 0.769(0.002)
N-GramXGB [50] 0.791(0.013) 0.691(0.008) 0.875(0.027) 0.655(0.007) 0.758(0.009) 0.787(0.004) 0.748(0.002)
PretrainGNN [51] 0.845(0.007) 0.687(0.013) 0.726(0.015) 0.627(0.008) 0.781(0.006) 0.799(0.007) 0.813(0.021)
GROVE_base [26] 0.821(0.007) 0.700(0.001) 0.812(0.030) 0.648(0.006) 0.743(0.001) 0.625(0.009) 0.673(0.018)
GROVE_large [26] 0.810(0.014) 0.695(0.001) 0.762(0.037) 0.654(0.001) 0.735(0.001) 0.682(0.011) 0.673(0.018)
GraphMVP [52] 0.812 (0.009) 0.724 (0.016) 0.791 (0.028) 0.639 (0.012) 0.759 (0.005) 0.770 (0.012) 0.777 (0.006)
MolCLR [53] 0.824 (0.009) 0.722 (0.021) 0.912 (0.035) 0.589 (0.014) 0.750 (0.002) 0.781 (0.005) 0.796 (0.019)
GEM [54] 0.856(0.011) 0.724(0.004) 0.901(0.013) 0.672(0.004) 0.781(0.001) 0.806(0.009) 0.817(0.005)
Mol-GDL [28] 0.863(0.019) 0.728(0.019) 0.966(0.002) 0.831(0.002) 0.794(0.005) 0.808(0.007) 0.675(0.014)
Uni-mol[55] 0.857 (0.002) 0.729 (0.006) 0.919 (0.018) 0.659 (0.013) 0.796 (0.005) 0.808 (0.003) 0.821 (0.013)
SMPT[56] 0.873 (0.015) 0.734 (0.003) 0.927 (0.002) 0.676 (0.050) 0.797 (0.001) 0.812 (0.001) 0.822 (0.008)
GNN-SKAN [38] 0.747(0.009) 0.676(0.014) - 0.614(0.005) 0.747(0.005) 0.786(0.015) -
KA-GCN 0.890(0.014) 0.787(0.014) 0.989(0.003) 0.842(0.001) 0.799(0.005) 0.821(0.005) 0.834(0.009)
KA-GAT 0.884(0.004) 0.785(0.021) 0.991(0.005) 0.847(0.002) 0.800(0.006) 0.823(0.002) 0.834(0.010)
Table 3: Comparison of the performance of KA-GNN (and KA-GAT) models with three types of base functions, including B-spline, polynomial, and Fourier series.
Dataset KA-GCN Models KA-GAT Models
B-spline Polynomial Fourier series B-spline Polynomial Fourier series
BACE 0.771(0.012) 0.853(0.027) 0.890(0.014) 0.808(0.009) 0.8319(0.007) 0.884(0.004)
BBBP 0.723(0.008) 0.654(0.009) 0.787(0.014) 0.657(0.004) 0.708(0.005) 0.785(0.021)
ChinTox 0.973(0.006) 0.981(0.004) 0.989(0.003) 0.948(0.003) 0.983(0.003) 0.991(0.005)
SIDER 0.824(0.003) 0.832(0.006) 0.842(0.001) 0.825(0.004) 0.836(0.001) 0.847(0.002)
Tox21 0.724(0.005) 0.715(0.004) 0.799(0.005) 0.731(0.012) 0.753(0.004) 0.800(0.006)
HIV 0.753(0.007) 0.804(0.010) 0.821(0.005) 0.744(0.018) 0.818(0.006) 0.823(0.002)
MUV 0.638(0.008) 0.787(0.012) 0.834(0.009) 0.792(0.013) 0.797(0.011) 0.834(0.010)
2.3.2 Importance of Fourier series for KA-GNNs
In this research, we employed a Fourier-based Kolmogorov-Arnold Network (KAN) as the primary computational mechanism within our GNN architecture, which differs significantly from traditional MLP-based approaches. To assess the impact of different base functions on the prediction of molecular properties, we explored several base function including B-spline, polynomial, Fourier-series, in the KA-GCNs architecture. These models were tested within the traditional GCN framework as well as the GAT framework. The comparative results are summarized in Table 3. The Polynomial-base function can be expressed in Eq (10).
$x_{j}^{(l+1)}=\sum_{i=1}^{n_{l}}\sum_{k=1}^{K}\left(C_{k,j,i}^{(l)}(x_{i}^{(l)})^{k}\right),$ (10)
where, $C_{k,j,i}^{(l)}$ is learnable parameters initially sampled from a normal distribution $\mathcal{N}\left(0,\frac{1}{n_{l+1}\times K}\right)$. $\mathbf{x}^{(l)}=(x_{1}^{(l)},x_{2}^{(l)},\dots,x_{n_{l}}^{(l)})$ denotes the input vector of activation values in layer $l$ of the Polynomial-series based KAN.
The data in the table illustrate that the KA-GCN and KA-GAT models, which incorporate Fourier-based methods, consistently outperform other models across all tested datasets. This underscores the substantial advantages of integrating Fourier-based approaches in enhancing the predictive accuracy and generalizability of GNN architectures. The superiority of these models is largely attributable to the powerful approximation capabilities of Fourier-based KAN, which we have substantiated in Theorem 2 and will further validate through subsequent function fitting evaluations. Furthermore, our ablation study not only validates our hypothesis that Fourier-based KAN provides a robust framework for improving molecular property prediction, significantly enhancing everything from molecular feature initialization to message passing in graph neural networks and prediction computations, but also compares the operational efficiency of models under different base functions. As shown in Figure 2 A and B, the results demonstrate that Fourier-based KAN significantly outperforms B-spline KAN in terms of efficiency. The Fourier-based models also clearly surpass the GCN and GAT models that utilize these other two base functions.
Table 4: Comparison of GCN/GAT models and KA-GCN/KA-GAT with Fourier series.
Dataset GCN KA-GCN GAT KA-GAT
BACE 0.835(0.014) 0.890(0.014) 0.834(0.012) 0.884(0.004)
BBBP 0.735(0.011) 0.787(0.014) 0.707(0.007) 0.785(0.021)
ChinTox 0.979(0.004) 0.989(0.003) 0.983(0.006) 0.991(0.005)
SIDER 0.834(0.001) 0.842(0.001) 0.836(0.002) 0.847(0.002)
Tox21 0.747(0.006) 0.799(0.005) 0.751(0.007) 0.800(0.006)
HIV 0.762(0.005) 0.821(0.005) 0.761(0.003) 0.823(0.002)
MUV 0.741(0.006) 0.834(0.009) 0.784(0.019) 0.834(0.010)
Furthermore, in this study, we also compared the traditional GCN and GAT models with our KA-GNNs (KA-GCN and KA-GAT), on the same feature molecular graphs. Table 4 illustrates that KA-GNNs demonstrate significant performance improvements across all datasets. This confirms that the incorporation of Fourier series significantly enhances the models’ ability to recognize and express the inherent periodicity and complexity in chemical structures.
Discussion
@@
We attribute the effectiveness and power of the KA-GNNs model to three core innovations. First, in constructing molecular graphs, we incorporate edges not only from traditional covalent bonds but also from non-covalent interactions based on a cut-off distance. This approach significantly enhances the model’s understanding of molecular structures, enabling it to capture more molecular properties than traditional covalent interactions alone. Second, we integrate the Kolmogorov-Arnold Network (KAN) into our model. KAN significantly reduces the number of parameters while providing high interpretability and enhanced expressive power, leading to superior performance of the KA-GNNs model. Third, we replace the pre-activation function in the KAN model from the original B-splines to Fourier series functions. Fourier series functions, widely used in signal processing, have proven effective in our model for analyzing and learning high-dimensional graphical data. By learning and optimizing the coefficients of these Fourier series functions for different variables, KA-GNNs improves the accuracy and stability of predictions when dealing with complex chemical and biological data structures. These three innovations collectively contribute to the outstanding performance of KA-GNNs, making it a powerful tool for molecular property prediction.
However, our KA-GNNs still have limitations. One of the key challenges is their interpretability. Even though KANs have been designed for excellent interpretability and indeed show good results in PDE examples. In our study, the pruning process however does not generate biologically meaningful superposition structures, that are easy to interpretate.
Figure 2: The comparison the computational efficiency of KA-GNNs based on B-spline, polynomial, and Fourier series. A. Running time of KA-GCN model for 100 epochs across different datasets. B. Running time of KA-GAT Model for 100 epochs across different datasets.
3 Methods
3.1 Approximation Capability of Fourier KAN
Given that the inner functions in Eq (1) of the Kolmogorov-Arnold representation theorem can exhibit significant non-smoothness, we do not use this theorem as the foundational theory for showing the approximation capability of our model. Instead, we base our approach on the extension of Carleson’s theorem [43] regarding the convergence of Fourier series for multivariable functions, as proved by Fefferman in [44]:
Theorem 3
[…]
almost everywhere.
Theorem 4
where $\circ$ denotes function composition.
@@ 3.1 Approximation Capability of Fourier KAN
Figure 3: Fourier-series KAN and two-layers MLP fit six different functions: A. Linear function $y=3x+5$, $K=200$ in Fourier-series KAN; B. Exponential function $y=3\exp(\frac{1}{2}x)$, $K=120$ in Fourier-series KAN; C. Logarithmic function $y=2\log(x)+3$, $K=100$ in Fourier-series KAN; D. Polynomial function $y=2x^{2}-4x+1$, $K=500$ in Fourier-series KAN; E. Sin function $y=2\sin(3x)+1$, $K=100$ in Fourier-series KAN; F. Sin and Cos function $y=3\sin(x)+2\cos(2x)+1$, $K=100$ in Fourier-series KAN.
In contrast to the multilayer perceptron (MLP) approach, which increases the number of layers to enhance the model’s ability to fit the data, the Fourier KAN model enhances its fitting ability by adjusting the number of Fourier basis functions (controlled by the K parameter) and the option to include a bias term to decompose complex functions into simpler sine and cosine waveforms. This strategic decomposition not only improves the model’s expressiveness but also reduces the computational burden. We evaluated the fitting ability of two-layers MLP and one-layer Fourier series KAN on six different function types, and the results are shown in Figure 3. Our analysis shows that Fourier KAN captures the underlying properties of the function more skillfully and exhibits superior expressiveness, especially in the ability to fit periodic functions.
3.2 Benchmark Datasets
Table 5: KA-GCN and KA-GAT parameters for different datasets.
Dataset Model Batch size LR K Layers Heads Epochs
BACE KA-GNN 128 1e-4 1 3 - 500
BACE KA-GAT 64 1e-4 2 2 4 500
BBBP KA-GNN 128 1e-4 2 1 - 500
BBBP KA-GAT 128 1e-4 5 2 2 500
ChinTox KA-GNN 128 1e-4 2 2 - 500
ChinTox KA-GAT 128 1e-4 5 2 2 500
SIDER KA-GNN 128 1e-4 2 1 - 500
SIDER KA-GAT 128 1e-4 4 2 2 500
Tox21 KA-GNN 512 1e-4 2 2 - 500
Tox21 KA-GAT 512 1e-4 3 2 2 500
HIV KA-GNN 512 1e-4 2 2 - 500
HIV KA-GAT 512 1e-4 2 2 4 500
MUV KA-GNN 512 1e-4 2 2 - 500
MUV KA-GAT 512 1e-4 3 2 4 500
3.3 Baseline Models
Our KA-GNN model has been evaluated alongside a selection of state-of-the-art GNN architectures, including both pre-trained and non-pre-trained models. These encompass MPNN frameworks like D-MPNN [49], attention-driven models such as AttentiveFP [48], and multi-scale approaches exemplified by Mol-GDL [28]. Importantly, our comparisons also extend to geometry-focused graph models that leverage pre-training, including N-Gram [50], PretrainGNN [51], GROVER [26], GraphMVP [52], MolCLR [53], GEM [54], Uni-mol [55], SMPT [56], and GNN-SKAN[38]. Table 5 is the hyperparameters setting for various datasets, including model type, batch size, learning rate (LR), number of harmonics ($K$), number of message passing layers, number of attention heads (for GAT), and number of epochs.
Table 6: Hyperparameter Sensitivity Analysis for KA-GNN.